A vessel with vertical sides and a base area of 0.0050 m2contains mercury of density 13 600 kg/m3and depth 0.60 m. Calculate

a. the weight of the mercury in the vessel.

b. the pressure that the mercury exerts on the base of the tank.
1
Expert's answer
2020-11-17T10:35:51-0500

Volume =area×depth

Volume=0.005m^2×0.6m

Volume=0.003m^3


Density=mass/volume

Mass=density×volume

Mass=13600×0.003

Mass=40.8kg


A. Weight=mass×g (g=acceleration due to gravity) g=10m/s^2

Weight=40.8×10=408N


B. Pressure =p×h×g (where p is density, h is height and g is acceleration due to gravity)

Pressure =13600×0.6×10

Pressure =81600mmhg.